# Hypergraph Foundation Model

### v1

@@ V Experiments
In this section, we collect and build ten text-attributed hypergraph datasets and conduct experiments on them to demonstrate the effectiveness of Hyper-FM.
@@ V-A Text-Attributed Hypergraph Datasets
This section introduces the text-attributed hypergraph datasets (TAHG) developed to support the Hypergraph Foundation Model. Recognizing that most existing hypergraph datasets feature vertex attributes primarily comprised of pre-extracted visual features or Bag-of-Words vector encodings, which are incompatible with large-scale models, we have manually constructed ten text-attributed hypergraph datasets. These datasets are categorized into four distinct classes to encompass a wide range of application domains.
The Citation Network Datasets encompass Cora-CA-Text, Cora-CC-Text, Pubmed-CA-Text, Pubmed-CC-Text, Aminer-Text, and Arxiv-Text. In these datasets, vertices represent papers with titles and abstracts serving as textual descriptions. For Cora and Pubmed, we collect textual descriptions based on the original paper IDs since this information is not available in current databases. Hyperedges are generated using co-author and co-citation relationships to capture the intricate interconnections within academic communities. For Arxiv-Text, we scrape data from the Arxiv repository, aggregating computer science papers published from 2010 to 2023 and selecting data vertices from five major categories: Computer Vision, Computation and Language, Robotics, Computers and Society, Cryptography and Security, and Software Engineering. Hyperedges are formed using co-citation relationships. Additionally, for Aminer-Text, we scrape data from the Aminer website and generate hyperedges based on co-author relationships.
@@
The statistics of these datasets are detailed in Tab. I. By incorporating rich textual attributes, these datasets enhance the capability of Hyper-FM to leverage semantic information across multiple domains, thereby facilitating more effective knowledge extraction and transfer learning in downstream tasks.
@@
Dataset #Classes #Text #V #E
Aminer-Text 3 Title+Abstract 8,226 4,050
Cora-CA-Text 7 Title+Abstract 2,708 1,922
Cora-CC-Text 7 Title+Abstract 2,708 2,161
Pubmed-CA-Text 3 Title+Abstract 19,717 28,154
Pubmed-CC-Text 3 Title+Abstract 19,717 10,547
Arxiv-Text 5 Title+Abstract 22,886 15,478
Movielens-Text 5 Name+Intro. 18,479 8,271
IMDB-Text 5 Name+Intro. 34,619 13,290
GoodBook-Text 4 Name+Detail 5,834 7,920
PPI-Text 3 Sequence 401 381
  • #V denotes “Number of vertices”, #E denotes “Number of hyperedges”, #Classes denotes “Number of classes”, #Text denotes “Type of text description”. “Intro.” is short for “Introduction”.
@@ V-B Experimental Settings
Our proposed method, Hyper-FM+Finetune, embodies the Hypergraph Foundation Model strategy. For hypergraph pretraining, we utilize two self-supervised techniques: HyperGCL [18], a representative contrastive learning framework tailored for hypergraphs, and SS-HT, which leverages the reconstruction of hypergraph vertex features as a self-supervised objective. The backbone of our foundation model incorporates both HGNN and HGNN+ architectures. In the HyperGCL+HGNN configuration, we apply the HyperGCL self-supervised pretraining strategy to the HGNN architecture. Subsequently, the pre-trained HGNN parameters are transferred and fine-tuned on downstream hypergraph datasets from different domains. This joint pretraining approach allows the model to integrate multi-domain structural knowledge, thereby enhancing its generalization capabilities across diverse downstream tasks.
@@
In our experiments, we utilize the bert-base-uncased[12] model as our language model for feature extraction. The neighborhood prediction task is configured with 4 layers, producing output feature dimensions of 768. To reduce complexity, we apply SVD to further compress the feature dimensions to 256. In terms of sampling strategy, we adopt a BFS sampling approach, selecting 500 vertices from each hypergraph domain and clustering them into 5 groups. We then generate hyperedges among the clustered vertices using a second-order K-Nearest Neighbors (KNN) method. In the fine-tuning phase, we adopt vertex classification as the downstream task. We employ a single-layer MLP as the classifier and fine-tune the pre-trained HGNN layers.
For the pre-training phase, we fix the number of HGNN layers at 2, with a hidden dimension of 128 for HGNN and HGNN+ models. Both the pre-training and fine-tuning phases utilize the Adam optimizer, with a learning rate of 0.001 during pre-training and a learning rate of 0.01 for the downstream vertex classification tasks. For a fair comparison, the hidden layer dimension for the supervised baseline (MLP, HGNN, HGNN+, HNHN and UniGCN) is also fixed at 128, learning late of Adam optimizer is also fixed at 0.01, matching the configurations of our main experiments. For fair comparisons, we apply the C-way-1-shot learning setting[47], for each dataset to build the training data and then allocate 100 vertices from each class as the validation set, while the remaining vertices are designated as the test set. We evaluate the performance of our model by computing the mean and variance of the results over five runs, providing a comprehensive understanding of the model’s effectiveness.
@@
Methods Aminer-Text Cora-CA-Text Cora-CC-Text Pubmed-CA-Text Pubmed-CC-Text
@@
  • The ”IMP” represents the improvement over the corresponding model without foundation model pretraining.
TABLE III: Experimental results on another five text-attributed hypergraph datasets.
Methods Arxiv-Text Movielens-Text IMDB-Text GoodBook-Text PPI-Text
Supervised MLP $0.3345_{\pm.02}$ $0.2880_{\pm.03}$ $0.2146_{\pm.09}$ $0.2489_{\pm.04}$ $0.4194_{\pm.09}$
HGNN $0.3523_{\pm.05}$ $0.3261_{\pm.08}$ $0.2533_{\pm.03}$ $0.3554_{\pm.03}$ $0.4306_{\pm.08}$
HGNN+ $0.3825_{\pm.08}$ $0.3016_{\pm.09}$ $0.2205_{\pm.09}$ $0.3225_{\pm.01}$ $0.5065_{\pm.08}$
HNHN $0.3753_{\pm.09}$ $0.2688_{\pm.09}$ $0.2528_{\pm.07}$ $0.3130_{\pm.01}$ $0.5261_{\pm.05}$
UniGCN $0.3571_{\pm.02}$ $0.3269_{\pm.07}$ $0.2565_{\pm.05}$ $0.3607_{\pm.04}$ $0.4387_{\pm.07}$
IP + Finetune HyperGCL + HGNN $0.3540_{\pm.06}$ $0.2384_{\pm.02}$ $0.2536_{\pm.03}$ $0.3278_{\pm.03}$ $0.4339_{\pm.07}$
HyperGCL + HGNN+ $0.3206_{\pm.04}$ $0.2796_{\pm.04}$ $0.2586_{\pm.04}$ $0.3422_{\pm.02}$ $0.5084_{\pm.06}$
SS-HT + HGNN $0.3505_{\pm.04}$ $0.3185_{\pm.05}$ $0.2486_{\pm.08}$ $0.3210_{\pm.01}$ $0.4355_{\pm.04}$
SS-HT + HGNN+ $0.3783_{\pm.07}$ $0.3174_{\pm.08}$ $0.2466_{\pm.08}$ $0.3166_{\pm.01}$ $0.4742_{\pm.03}$
Hyper-FM + Finetune HyperGCL + HGNN $0.4189_{\pm.04}$ $0.3464_{\pm.08}$ $0.2676_{\pm.04}$ $0.3588_{\pm.04}$ $0.5371_{\pm.08}$
HyperGCL + HGNN+ $0.4198_{\pm.07}$ $\mathbf{0.3470_{\pm.06}}$ $0.2672_{\pm.04}$ $\mathbf{0.3703_{\pm.04}}$ $\mathbf{0.5742_{\pm.06}}$
SS-HT + HGNN $0.4044_{\pm.05}$ $0.3305_{\pm.05}$ $\mathbf{0.2786_{\pm.08}}$ $0.3589_{\pm.04}$ $0.5403_{\pm.04}$
SS-HT + HGNN+ $\mathbf{0.4649_{\pm.05}}$ $0.3347_{\pm.09}$ $0.2775_{\pm.08}$ $0.3569_{\pm.01}$ $0.5387_{\pm.03}$
IMP(%) $\mathbf{21.5\%}$ $\mathbf{15.1\%}$ $\mathbf{9.9\%}$ $\mathbf{14.8\%}$ $\mathbf{13.4\%}$
@@ V-C Results and Discussions
Tabs. II and III present the experimental results of our proposed Hyper-FM on ten TAHG datasets. From these tables, we can draw the following four key observations.
Firstly, hypergraph neural network methods such as HGNN consistently outperform the MLP baseline across all datasets. This improvement underscores the effectiveness of leveraging hypergraph-structured associations, which facilitate the transmission of relational information and enable the model to capture more comprehensive features, thereby enhancing overall performance. Secondly, our proposed Hyper-FM method demonstrates a substantial performance enhancement of approximately 13.3% across all ten datasets, with an exceptional improvement of 23.4% observed on the Pubmed-CC-Text dataset. This significant uplift highlights the efficacy of the Hypergraph Foundation Model. By conducting pretraining on multi-domain hypergraph datasets, Hyper-FM successfully integrates diverse domain-specific knowledge into the hypergraph, which in turn markedly boosts the performance of target domain models during downstream tasks. Thirdly, when comparing with pretraining-based methods, the IP+Finetune approach exhibits noticeably inferior performance relative to Hyper-FM+Finetune. In particular, the Cora-CC-Text dataset experiences negative transfer, resulting in decreased performance in the target domain. This decline can be attributed to the sequential nature of IP+Finetune’s pretraining strategy, which, despite utilizing self-supervised training to acquire multi-modal hypergraph domain knowledge, is susceptible to knowledge forgetting. Additionally, the significant disparities in knowledge distributions across different domains hinder the convergence of sequential pretraining, thereby adversely affecting performance on downstream tasks.
Lastly, within the Hypergraph Foundation Model framework, the choice of self-supervised strategy—whether HyperGCL or SS-HT—has a relatively minor impact on the overall results. In contrast, the utilization of multi-domain hierarchical hypergraphs exerts a pronounced effect on performance. A similar trend is observed in the graph foundation model GCOPE[47], reinforcing the conclusion that the construction of robust association structures and the effective integration of cross-domain association information are critical focal points for advancing hypergraph foundation models. These observations collectively demonstrate that our Hyper-FM not only leverages the inherent advantages of hypergraph structures but also effectively integrates multi-domain knowledge through sophisticated pretraining strategies. This dual capability underpins the model’s superior performance and highlights the importance of comprehensive structural and semantic knowledge integration in hypergraph-based learning frameworks.
@@ V-D More Evaluations and Analysis
In this subsection, we conduct ablation studies to evaluate the individual contributions of the Vertex Knowledge Embedding module and the Structural Knowledge Extraction module within our Hyper-FM on four TAHG datasets.
@@ V-D1 Ablation Study on Vertex Knowledge Embedding
Tab. IV presents the experimental results for various vertex feature extraction methods. Specifically, we compare the performance of Bag-of-Words (BoW), direct feature extraction using the LM, and our proposed approach, LM with Neighborhood Prediction fine-tuning (LM+NP). The BoW method serves as a naive baseline by representing vertex features based on word frequency, while the LM approach leverages pre-trained language models to extract semantic features directly from the textual data. The LM+NP method enhances the LM by incorporating neighborhood information through a Neighborhood Prediction task during fine-tuning. From the results, it is evident that there is no significant difference in performance between the LM and BoW methods overall. Notably, the LM features demonstrate superior performance on the Pubmed-CA-Text and Movielens-Text datasets, whereas the BoW features yield better results on the GoodBook-Text and PPI-Text datasets. This observation indicates that the BoW approach, despite its simplicity, is capable of capturing essential vertex feature information in certain contexts. However, when employing the proposed LM+NP method, we observe a substantial improvement in performance across all four datasets. Specifically, the introduction of neighborhood knowledge through fine-tuning enables the model to generate more informative and rich vertex embeddings, leading to enhanced accuracy in downstream tasks. These findings underscore the advantage of integrating neighborhood information into vertex feature representations. While both BoW and LM methods have their respective strengths, the LM+NP approach effectively combines the semantic richness of language models with neighborhood insights, resulting in more robust and accurate vertex embeddings. Consequently, the enhanced vertex knowledge embedded through our proposed method plays a crucial role in improving the overall performance of the Hypergraph Foundation Model on diverse Text-Attributed Hypergraph datasets.

### v2

@@ V Experiments
In this section, we collect and build eleven text-attributed hypergraph datasets and conduct experiments on them to demonstrate the effectiveness of Hyper-FM.
@@ V-A Text-Attributed Hypergraph Datasets
This section introduces the text-attributed hypergraph datasets (TAHG) developed to support the Hypergraph Foundation Model. Recognizing that most existing hypergraph datasets feature vertex attributes primarily comprised of pre-extracted visual features or Bag-of-Words vector encodings, which are incompatible with large-scale models, we have manually constructed eleven text-attributed hypergraph datasets. These datasets are categorized into four distinct classes to encompass a wide range of application domains.
The Citation Network Datasets encompass Cora-CA-Text, Cora-CC-Text, Pubmed-CA-Text, Pubmed-CC-Text, Aminer-Text, Arxiv-Text, and OGBN-Arxiv. In these datasets, vertices represent papers with titles and abstracts serving as textual descriptions. For Cora and Pubmed, we collect textual descriptions based on the original paper IDs since this information is not available in current databases; Hyperedges are generated using co-author and co-citation relationships to capture the intricate interconnections within academic communities. For Arxiv-Text, we scrape data from the Arxiv repository, aggregating computer science papers published from 2010 to 2023, selecting data vertices from five major categories (Computer Vision, Computation and Language, Robotics, Computers and Society, Cryptography and Security, Software Engineering), and forming hyperedges via co-citation relationships. Additionally, for Aminer-Text, we scrape data from the Aminer website and generate hyperedges based on co-author relationships. OGBN-Arxiv is an additional large-scale dataset directly adopted from public resources [^1] https://ogb.stanford.edu/docs/nodeprop/. It is a directed citation network of Computer Science arXiv papers indexed by MAG [54], where each vertex denotes an arXiv paper with textual descriptions consisting of the original paper’s title and abstract. Each directed edge represents a citation relationship, and its task is 40-class subject area prediction (e.g., cs.AI, cs.LG).
@@
The statistics of these datasets are given in Tab. I. By incorporating rich textual attributes, these datasets enhance the capability of Hyper-FM to leverage semantic information across multiple domains, thereby facilitating more effective knowledge extraction and transfer learning in downstream tasks.
@@
Dataset #Label #Text #Avg. #V #E
Aminer-Text 3 Title+Abstract 215.45 8,226 4,050
Cora-CA-Text 7 Title+Abstract 155.89 2,708 1,922
Cora-$C$-Text 7 Title+Abstract 155.89 2,708 2,161
Pubmed-CA-Text 3 Title+Abstract 249.16 19,717 28,154
Pubmed-$C$-Text 3 Title+Abstract 249.16 19,717 10,547
Arxiv-Text 5 Title+Abstract 194.10 22,886 15,478
Movielens-Text 5 Name+Intro. 43.26 18,479 8,271
IMDB-Text 5 Name+Intro. 57.43 34,619 13,290
GoodBook-Text 4 Name+Detail 182.53 5,834 7,920
PPI-Text 3 Sequence 587.84 401 381
OGBN-Arxiv 40 Title+Abstract 170.36 169,343 168,537
  • #V denotes “Number of vertices”, #E denotes “Number of hyperedges”, #Classes denotes “Number of classes”, #Text denotes “Type of text description”, #Avg. denotes “Average length of vertex text descriptions”. “Intro.” is short for “Introduction”.
@@ V-B Experimental Settings
Our proposed method, Hyper-FM+Finetune, embodies the Hypergraph Foundation Model strategy. For hypergraph pretraining, we utilize two self-supervised techniques: HyperGCL [18], a representative contrastive learning framework tailored for hypergraphs, and SS-HT[51], which leverages the reconstruction of hypergraph vertex features as a self-supervised objective. The backbone of our foundation model incorporates both HGNN and HGNN+ architectures. In the HyperGCL+HGNN configuration, we apply the HyperGCL self-supervised pretraining strategy to the HGNN architecture. Subsequently, the pre-trained HGNN parameters are transferred and fine-tuned on downstream hypergraph datasets from different domains. This joint pretraining approach allows the model to integrate multi-domain structural knowledge, thereby enhancing its generalization capabilities across diverse downstream tasks.
@@
We first conduct data preprocessing, including hyperedge construction, hypergraph structure optimization and vertex feature processing. When building hyperedges, any co-citation or co-director-like association forms a hyperedge; however, during dataset construction, we remove hyperedges with abnormally large degrees (>40>40, typically accounting for a small proportion) to reduce structural noise, and further eliminate isolated vertices from the remaining hypergraph to ensure valid node connectivity. For vertex features, we use the bert-base-uncased[12] model to extract 768-dimensional textual features, then compress the dimension to 256 via Singular Value Decomposition (SVD) to lower computational complexity.
In the pretraining phase, we fix HGNN layers at 2, with a hidden dimension of 128 for both HGNN and HGNN+. The Vertex Knowledge Embedding module uses 4 hierarchical layers ($T=4$, validated in [55]); the number of hierarchical label clusters $K$ is calculated as $K=N/B$ ($B=100$, NN is sample count), following [47]—excessively small/large $K$ degrades performance, especially small $K$ leading to overcrowded clusters. While we adopt this fixed heuristic for consistency with prior work, we acknowledge that adaptive or dynamic determination of $K$ within constrained ranges could further enhance model adaptability and performance, and we leave this as an important direction for future exploration. For ”Multi-Domain Hypergraphs Sampling”, we adopt BFS: neighbor vertices are randomly sampled at $80\%\sim 100\%$, expansion depth is constrained by total sampled vertices, and the sampling rate is dynamically adjusted to cover at least third-order neighbors of each vertex. After sampling 500 vertices per domain, we cluster them into 5 groups and generate hyperedges via second-order KNN. The Adam optimizer with a learning rate of 0.001 is used here.
In the fine-tuning phase, vertex classification serves as the downstream task, with a single-layer MLP as the classifier and only pre-trained HGNN layers fine-tuned to retain generalizable pretraining knowledge. For fair comparison, supervised baselines (MLP, HGNN, HGNN+, HNHN, UniGCN) have a hidden dimension of 128 (matching the pretraining backbone) and use Adam with a learning rate of 0.01 (consistent with the downstream task). Data splitting follows the C-way-1-shot setting[53]: training data is built under this setting, 100 vertices per class are used as the validation set, and the rest as the test set. For datasets with temporal information (e.g., OGBN-Arxiv), all samples were shuffled across publication years before random splitting, since chronological order is not the focus of our study and random sampling avoids potential bias. Performance is evaluated by the mean and variance of five independent runs to ensure reliability.
@@
Methods Aminer-Text Cora-CA-Text Cora-$C$-Text Pubmed-CA-Text Pubmed-$C$-Text
@@
  • The “IMP” denotes the performance gain of the optimal method under the “Hyper-FM + Finetune” setting relative to its counterpart without pretraining.
TABLE III: Experimental results on another six text-attributed hypergraph datasets.
Methods Arxiv-Text Movielens-Text IMDB-Text GoodBook-Text PPI-Text OGBN-Arxiv
Supervised MLP $0.3345_{\pm.02}$ $0.2880_{\pm.03}$ $0.2146_{\pm.09}$ $0.2489_{\pm.04}$ $0.4194_{\pm.09}$ $0.2034_{\pm.04}$
HGNN $0.3523_{\pm.05}$ $0.3261_{\pm.08}$ $0.2533_{\pm.03}$ $0.3554_{\pm.03}$ $0.4306_{\pm.08}$ $0.2348_{\pm.04}$
HGNN+ $0.3825_{\pm.08}$ $0.3016_{\pm.09}$ $0.2205_{\pm.09}$ $0.3225_{\pm.01}$ $0.5065_{\pm.08}$ $0.2441_{\pm.03}$
HNHN $0.3753_{\pm.09}$ $0.2688_{\pm.09}$ $0.2528_{\pm.07}$ $0.3130_{\pm.01}$ $0.5261_{\pm.05}$ $0.2331_{\pm.03}$
UniGCN $0.3571_{\pm.02}$ $0.3269_{\pm.07}$ $0.2565_{\pm.05}$ $0.3607_{\pm.04}$ $0.4387_{\pm.07}$ $0.2559_{\pm.04}$
IP + Finetune HyperGCL + HGNN $0.3540_{\pm.06}$ $0.2384_{\pm.02}$ $0.2536_{\pm.03}$ $0.3278_{\pm.03}$ $0.4339_{\pm.07}$ $0.2657_{\pm.04}$
HyperGCL + HGNN+ $0.3206_{\pm.04}$ $0.2796_{\pm.04}$ $0.2586_{\pm.04}$ $0.3422_{\pm.02}$ $0.5084_{\pm.06}$ $0.2665_{\pm.04}$
SS-HT + HGNN $0.3505_{\pm.04}$ $0.3185_{\pm.05}$ $0.2486_{\pm.08}$ $0.3210_{\pm.01}$ $0.4355_{\pm.04}$ $0.2414_{\pm.04}$
SS-HT + HGNN+ $0.3783_{\pm.07}$ $0.3174_{\pm.08}$ $0.2466_{\pm.08}$ $0.3166_{\pm.01}$ $0.4742_{\pm.03}$ $0.2596_{\pm.04}$
Hyper-FM + Finetune HyperGCL + HGNN $0.4189_{\pm.04}$ $0.3464_{\pm.08}$ $0.2676_{\pm.04}$ $0.3588_{\pm.04}$ $0.5371_{\pm.08}$ $0.2764_{\pm.01}$
HyperGCL + HGNN+ $0.4198_{\pm.07}$ $\mathbf{0.3470_{\pm.06}}$ $0.2672_{\pm.04}$ $\mathbf{0.3703_{\pm.04}}$ $\mathbf{0.5742_{\pm.06}}$ $\mathbf{0.2789_{\pm.01}}$
SS-HT + HGNN $0.4044_{\pm.05}$ $0.3305_{\pm.05}$ $\mathbf{0.2786_{\pm.08}}$ $0.3589_{\pm.04}$ $0.5403_{\pm.04}$ $0.2759_{\pm.04}$
SS-HT + HGNN+ $\mathbf{0.4649_{\pm.05}}$ $0.3347_{\pm.09}$ $0.2775_{\pm.08}$ $0.3569_{\pm.01}$ $0.5387_{\pm.03}$ $0.2745_{\pm.04}$
IMP(%) $\mathbf{21.5\%}$ $\mathbf{15.1\%}$ $\mathbf{9.9\%}$ $\mathbf{14.8\%}$ $\mathbf{13.4\%}$ $\mathbf{14.3\%}$
@@ V-C Results and Discussions
Tabs. II and III present the experimental results of our proposed Hyper-FM on eleven TAHG datasets. From these tables, we can draw the following four key observations.
Firstly, hypergraph neural network methods such as HGNN consistently outperform the MLP baseline across all datasets. This improvement underscores the effectiveness of leveraging hypergraph-structured associations, which facilitate the transmission of relational information and enable the model to capture more comprehensive features, thereby enhancing overall performance. Secondly, our proposed Hyper-FM method demonstrates a substantial performance enhancement of approximately 13.4% across all eleven datasets, with an exceptional improvement of 23.4% observed on the Pubmed-CC-Text dataset. Notably, we also observe a significant improvement of 14.3% on the large-scale and complex vertex classification dataset OGBN-Arxiv, which demonstrates the effectiveness of Hyper-FM in handling large-scale and complex hypergraph data. This significant uplift highlights the efficacy of the Hypergraph Foundation Model. By conducting pretraining on multi-domain hypergraph datasets, Hyper-FM successfully integrates diverse domain-specific knowledge into the hypergraph, which in turn markedly boosts the performance of target domain models during downstream tasks. Thirdly, when comparing with pretraining-based methods, the IP+Finetune approach exhibits noticeably inferior performance relative to Hyper-FM+Finetune. In particular, the Cora-CC-Text dataset experiences negative transfer, resulting in decreased performance in the target domain. This decline can be attributed to the sequential nature of IP+Finetune’s pretraining strategy, which despite utilizing self-supervised training to acquire multi-modal hypergraph domain knowledge, is susceptible to knowledge forgetting. Additionally, the significant disparities in knowledge distributions across different domains hinder the convergence of sequential pretraining, thereby adversely affecting performance on downstream tasks.
Lastly, within the Hypergraph Foundation Model framework, the choice of self-supervised strategy, whether HyperGCL or SS-HT, has a relatively minor impact on the overall results. In contrast, the utilization of multi-domain hierarchical hypergraphs exerts a pronounced effect on performance. A similar trend is observed in the graph foundation model GCOPE[53], reinforcing the conclusion that the construction of robust association structures and the effective integration of cross-domain association information are critical focal points for advancing hypergraph foundation models. These observations collectively demonstrate that our Hyper-FM not only leverages the inherent advantages of hypergraph structures but also effectively integrates multi-domain knowledge through sophisticated pretraining strategies. This dual capability underpins the model’s superior performance and highlights the importance of comprehensive structural and semantic knowledge integration in hypergraph-based learning frameworks.
@@ V-D More Evaluations and Analysis
In this subsection, we conduct ablation studies to evaluate the individual contributions of the Vertex Knowledge Embedding module and the Structural Knowledge Extraction module within our Hyper-FM on eleven TAHG datasets.
@@ V-D1 Ablation Study on Vertex Knowledge Embedding
Tab. IV presents the experimental results for various vertex feature extraction methods. Specifically, we compare the performance of Bag-of-Words (BoW), direct feature extraction using the LM, and our proposed approach, LM with Neighborhood Prediction fine-tuning (LM+NP). The BoW method serves as a naive baseline by representing vertex features based on word frequency, while the LM approach leverages pre-trained language models to extract semantic features directly from the textual data. The LM+NP method enhances the LM by incorporating neighborhood information through a Neighborhood Prediction task during fine-tuning. From the results, it is evident that there is no significant difference in performance between the LM and BoW methods overall. Notably, the LM features demonstrate superior performance on the Pubmed-CA-Text, Movielens-Text, Aminer-Text, Cora-CA-Text, and Cora-CC-Text datasets, whereas the BoW features yield better results on the GoodBook-Text and PPI-Text datasets. This observation indicates that the BoW approach, despite its simplicity, is capable of capturing essential vertex feature information in certain contexts. However, when employing the proposed LM+NP method, we observe a substantial improvement in performance across all eleven datasets. Specifically, the introduction of neighborhood knowledge through fine-tuning enables the model to generate more informative and rich vertex embeddings, leading to enhanced accuracy in downstream tasks. These findings underscore the advantage of integrating neighborhood information into vertex feature representations. While both BoW and LM methods have their respective strengths, the LM+NP approach effectively combines the semantic richness of language models with neighborhood insights, resulting in more robust and accurate vertex embeddings. Consequently, the enhanced vertex knowledge embedded through our proposed method plays a crucial role in improving the overall performance of the Hypergraph Foundation Model on diverse Text-Attributed Hypergraph datasets.
@@
Aminer-Text $0.3940_{\pm.01}$ $0.4073_{\pm.01}$ $\mathbf{0.4325_{\pm.02}}$
Cora-CA-Text $0.2101_{\pm.02}$ $0.2526_{\pm.05}$ $\mathbf{0.2633_{\pm.03}}$
Cora-$C$-Text $0.3059_{\pm.04}$ $0.3368_{\pm.04}$ $\mathbf{0.3706_{\pm.04}}$
Pubmed-$C$-Text $0.3977_{\pm.01}$ $0.3967_{\pm.01}$ $\mathbf{0.4009_{\pm.01}}$
Arxiv-Text $0.2786_{\pm.03}$ $0.3231_{\pm.07}$ $\mathbf{0.4189_{\pm.04}}$
IMDB-Text $0.2625_{\pm.05}$ $0.2637_{\pm.04}$ $\mathbf{0.2676_{\pm.04}}$
OGBN-Arxiv $0.2667_{\pm.02}$ $0.2608_{\pm.03}$ $\mathbf{0.2789_{\pm.01}}$